What is the change in surface energy, when a mercury drop of radius R splits up into 1000 droplets of radius r?
(a) 8 pR2T (b) 16 pR2T
(c) 24 pR2T (d) 36 pR2T

Grade:12

2 Answers

Arun
3 years ago

Surface Energy -

It is defined as the amount of work done in increasing the area of the liquid against surface tension.

-

 

  n\frac{4}{3}\pi r^{3}=\frac{4}{3}\pi R^{3}

nr^{3}=R^{3}

=> R=n^{1/3}r

where n= 10000

After calculeting we get r=36\rho R^2T

Vikas TU
3 years ago
Surface Energy -
It is defined as the amount of work done in increasing the area of the liquid against surface tension.
where n= 10000
After calculeting we get r=36\rho R^2T